OpenStudy (anonymous):

What is the solution of the following system? {3x-3y=10 -9x-9y=-30

OpenStudy (paounn):

multiply 1st equation by 3 and sum to the second, you get \[-18y = 0 \rightarrow y=0\] by substituting thatresult in any of the equation (easier in the first one) you get \[3x -3*0 = 10 \rightarrow 3x = 10 x= 10/3\]. solution is then \[x={10\over 3}; y=0\]

hero (hero):

@Burningitdown double-check for sign errors

OpenStudy (anonymous):

Uhm. @Hero I think there may have been one, yeah.

OpenStudy (anonymous):

@Paounn Thanks. I made a mistake in typing the problem, but I appreciate your effort.

OpenStudy (anonymous):

It should have been, {3x+3y=10 -9x-9y=-30

hero (hero):

Now that we have corrected the mistakes, Divide both sides of the second equation by -3 to get 3x + 3y = 10 Then realize that both equations are the same. Thus infinite solutions

OpenStudy (paounn):

Now that it's fixed ( :D ) you can see that second equation is obtained from the first by multiplying by -3 b oth terms: they are, in fact, the *same* equation. System admits \[\infty^1\] solutions, given by \[x = t, y= {{10-3t} \over 3} t \mathbb{R}\]

OpenStudy (anonymous):

You both basically said the same thing, so I'll fan you both. Thanks.

OpenStudy (anonymous):

I greatly appreciate it.

hero (hero):

But without me, you wouldn't have found your mistake.